# Can Large Language Models Improve Phishing Defense? A Large-Scale Controlled Experiment on Warning Dialogue Explanations

### v1

@@ 3 Approach
Regarding the explanation style, the first approach we considered is based on feature-based explanations, a widely adopted approach in HCI systems for communicating the most influential features contributing to an AI decision (Lai et al., 2023; Fok and Weld, 2024). In our context, these explanations highlight specific elements of an email, such as a masked link or a suspicious sender address, that led the AI to classify it as phishing. Complementing this standard, we also include counterfactual explanations, which offer an alternative, theoretically grounded approach to supporting user understanding. Rather than simply highlighting influential features, counterfactuals offer contrastive explanations answering implicit “why not?” questions by describing how small, meaningful changes would alter the AI’s prediction (Lee and Chew, 2023; Koh et al., 2024; Gentile et al., 2025). This explanation technique clarifies decision boundaries and supports actionable recourse, particularly relevant in high-stakes settings such as fraud detection or security. In our phishing warning dialogues, we embed these explanations to show how the email could have avoided being flagged, for example, stating that the email would be considered safe if the link label matched the actual destination URL (e.g., “protect your account” linking to facebook.com/protect-account).
The second factor concerns the LLM model used to generate the explanation message. While past studies often relied on manually written explanation messages demonstrating their effectiveness (Desolda et al., 2023; Greco et al., 2025; Buono et al., 2023), the growing availability of LLMs might enable automatic generation of tailored content in both explanation styles. If explanations produced by LLMs prove as effective as manually crafted messages, which are both time-consuming to create and must be updated whenever new features are introduced, then LLMs could fully automate their generation. This study considered explanations generated by an open-source model (Llama 3.3) and a closed-source commercial model (Claude 3.5 Sonnet). These two LLMs were considered to evaluate whether a commercially available LLM (Claude) could provide better explanations than an open-weight solution such as LLama 3.3. This aspect is critical, as an organization could have strict privacy requirements that prohibit the sharing of sensitive data (i.e., emails) with third parties like Claude (Shashidhar et al., 2023; Panwar et al., 2024; Zhang et al., 2024; Adams et al., 2024; López Espejel et al., 2025).
@@
  • 4.
    RQ4: How do user-related factors affect click trough rate (CRT) of phishing links?
@@ 3.1 Warning dialogues design
The phishing warning dialogues explored in this study have the same structure to isolate the effects of the explanation message, depicted in Figure 1. A dialogue opens with an attention-grabbing header flagging the link as potentially dangerous, followed by a summary that attributes this judgment to the AI system. The main part of the interface is the explanation message, presented as a short paragraph embedded within the dialogue box. The bottom part contains buttons to get back to the email visualization or, with a less prominent presentation, a link to continue browsing the destination. All messages have the same typographic style and layout to control for visual biases. This design ensured that any differences in user response could be attributed to the nature of the explanations rather than variations in formatting.
Figure 1: Anatomy of the warning dialogue used in the study. The top part (A) includes a brief warning working as the header of the box, the main part (B) presents the explanation message, while the footer (C) shows controls to get back to the email visualization or to proceed with browsing.
@@ 3.2 Explanation Styles
The explanation message included in the main part differs according to the considered styles. Feature-based explanations emphasize observable cues extracted from the email, such as sender identity mismatches or suspicious links. The left part of Figure 4 shows sample explanations for this style, realised through declarative statements pointing to concrete phishing indicators. All messages share a common structure consisting of three sentences, inspired by prior work Desolda et al. (2023), which builds on the Communication-Human Information Processing (C-HIP) model (Wogalter, 2018) and previous studies on effective warning communication (Bauer et al., 2013). The same structure was explicitly adopted when prompting the LLMs to generate the warnings, ensuring consistency across human- and AI-authored messages.
In contrast, counterfactual explanations present a hypothetical scenario in which the email would not have been flagged. These explanations reframe the AI’s decision as contingent on a small change, inviting users to consider alternatives. The examples adopt a conditional format, underscoring a key difference in framing: while feature-based explanations justify the current decision, counterfactuals highlight what could have made the outcome different. The right part of Figure 4 shows sample warnings using a counterfactual style.
@@ 3.3 Generating the explanations
The other difference in the experimental conditions is the LLM used to generate the warning text. We selected the two LLMs to create explanations for this study through a rigorous evaluation process. Notably, 13 different LLMs (six commercially available, seven open-source) were included in our evaluation process as possible candidates. As grounds for this evaluation, we considered the three emails (two true positives and one false positive) whose links, when clicked, generated a warning dialogue during the experiment. To evaluate each LLM, we followed this process:
  • 1.
    Following the approach and the prompts proposed in Desolda et al. (2025), for each of the three emails, the LLM was prompted to generate an explanation message to be included in the warning, given the email body and one of the features to explain. We considered the following three email features, which were found to be helpful to users when making decisions regarding phishing content (Greco et al., 2025; Desolda et al., 2025), i.e., (i) the Top-Level Domain of the URL in the email is misplaced (i.e., “www.amazon.com.cz”); (ii) the URL in the email is an IP address; (iii) the link shown in the email and its actual destination mismatch (e.g., ”click here” is shown instead of ”www.facebook.com/…”) — this feature was found in the false positive email.
  • 2.
    For each of the three emails, we generated a counterfactual explanation message through a modified version of the prompt proposed in Desolda et al. (2025).
  • 3.
@@
  • 4.
    The explanations underwent an automatic measurement of text readability with the FK (Flesch-Kincaid) Grade Level (Kincaid, 1975) and of text understandability with the SMOG (Simple Measure of Gobbledygook) Index (Laughlin, 1969).
  • 5.
@@
To select the best commercial and open-weight LLMs, we initially selected a set of LLMs with the lowest, thus best, SMOG Index and FK Grade for both feature-based and counterfactual explanations. From this analysis, we selected Claude 3.5 Sonnet and Gemini 1.5 Flash as the best candidates for commercial LLMs, while all the Llama LLMs are candidates for the open-weight solutions. Then, two experts in HCI and cybersecurity performed a manual inspection of the generated explanations. This analysis removed Gemini 1.5 Flash because it generated explanations that were either too generic or contained hallucinations, especially for the counterfactual style (e.g., for the Amazon phishing email it generated an explanation reporting “[..] A safe email address wouldn’t have strange parts in its name. [..]”). Moreover, Llama 3.2 90B (as well as Gemini 1.5 Pro) was excluded, as it could not be forced to produce an explanation for the false-positive email. Llama 3.1 8B was a good candidate for feature-based explanations, although readability and understandability metrics were mediocre in the context of counterfactual explanations. Moreover, the generated feature-based explanation for the Amazon email contained a hallucination (“[..] It looks like an Amazon link, but the ’.com’ name is typically found in other types of companies. [..]”). Therefore, considering the qualitative assessment and the average metrics for both explanation styles, we selected Llama 3.3 70B as it represents a good trade-off between the resources needed to deploy this LLM and the quality of the explanations. The explanations produced by Claude 3.5 Sonnet and Llama 3.3 70B were then featured in the user study as described in Section 4.1.
Figure 2: Average understandability and readability metrics for the feature-based explanations generated by the different LLMs - the lower, the better. The asterisks indicate the LLMs that were finally chosen.
Figure 3: Average understandability and readability metrics for the counterfactual explanations generated by the different LLMs - the lower, the better. The asterisks indicate the LLMs that were finally chosen.
@@ 3.4 Claude vs Llama
Instead, Llama delivers explanations that are somewhat more concise and technical in tone. Its feature-based outputs highlight specific anomalies (e.g., numeric IP addresses, suspicious domain extensions) with minimal elaboration. The counterfactuals follow a similar structure to Claude’s but are generally more direct and less descriptive, focusing on the change required to alter the classification without extending into consequences. Llama also tends to use simpler language and more neutral phrasing, which may be more mechanical but still effective. These stylistic differences are illustrated in Figure 4. The top row includes examples generated with Claude 3.5 Sonnet, while the bottom row includes examples generated with Llama 3.0 70B.
Figure 4: Four versions of the warning dialogue for the same deceptive link. Rows correspond to the text generated by the two LLMs employed in this study (Claude 3.5 Sonnet and Llama 3.0 70B), while columns refer to explanation styles (feature-based and counterfactual).
@@ 4 Methods
In this section, we present a controlled experiment that compares the effectiveness of various explanations displayed in warning dialogues to protect users from phishing emails. The study specifically focused on evaluating explanations based on two factors we identified and discussed in the previous section: the explanation style (either feature-based or counterfactual) and the LLM model used to generate the explanation message (LLama 3.3 or Claude 3.5 Sonnet). We also considered a baseline warning including manually-generated messages that was empirically demonstrated to outperform all the existing warnings for phishing attacks (Greco et al., 2025). The study procedure and materials (i.e., warning design and emails included) were identical to the study in Greco et al. (2025) to make the comparison with the baseline as fair and rigorous as possible.
@@ 4.1 Study Design
This study adopted a between-subjects design with the explanation style and LLM used to generate the explanations as the independent variables. Five total between-subject levels were considered, four experimental warnings plus a baseline:
@@
    Baseline: warnings that included feature-based explanations, which were manually written by experts (existing dataset - from Greco et al. (2025))
To improve the external validity of the study, for each experimental condition, we generated three explanations related to three different features. The complete list of explanation messages contained in each warning is reported in A.
@@ 4.2 Participants
A total of 750 participants were involved, with equal allocation (n = 150) to each condition. The manually generated feature-based explanation condition included existing data (collected separately, n = 150), while the remaining 600 participants were recruited specifically for this study through the Prolific online platform [^2] https://www.prolific.co. The inclusion criteria for participants were as follows: i) fluency in English, ii) an even split between men and women (50-50), and iii) 18 or more years of age. A total of 61 participants either failed attention checks or were identified as inattentive and were thus excluded. To balance participation across the experimental conditions, we recruited 61 more participants to reach a final sample of exactly 750 participants. The final sample had an average age of 32.98 years (sd = 11.01) and reported spending 7.94 hours per day on the Internet (sd = 3.95). Participants were 375 males and 375 females. The participation in the study lasted approximately 20 minutes, and participants were rewarded with £3.00, in line with Prolific’s recommended participation fee of £9.00/hour.
@@ 4.3 Materials
A warning dialogue, such as that shown in Figure 1, was shown to users who clicked on a phishing link. To simulate the case in which the system mistakenly classifies a genuine email as phishing, i.e., a false positive situation, we forced the platform to show a warning also for one of the genuine emails. After the study, each user may have seen a maximum of three warning dialogues.
A full suite of user interactions was recorded in this study: hovering over phishing links, clicking on them, visiting associated URLs, and responding to warnings, which allowed calculating the click-through rate (CTR) for each participant. The CTR was the proportion of accessed suspicious links to the number of displayed warnings, and it was an objective measure of the effectiveness of the warning.
The survey instrument employed by Desolda et al. (2023) was administered to comprehend users’ perceptions of the warnings. All the questions and their formats of responses are given in Table 1.
To test for the effects of users’ knowledge about basic cybersecurity concepts, a 10-item test questionnaire was used (Olmstead and Smith, 2017).
Finally, a short, 6-item version of the Need for Cognition Scale (NCS-6) (Coelho et al., 2020) was used to measure users’ Need for Cognition, which is the tendency to engage in and enjoy effortful cognitive activities (Cacioppo and Petty, 1982). Previous work highlighted that individual differences like NFC might significantly impact the effectiveness of AI support and explanations in terms of performance, overreliance on AI, and learning (Cacioppo et al., 1996; Buçinca et al., 2021; Gajos and Mamykina, 2022; Buçinca et al., 2024)
Table 1: The questionnaire used to evaluate the warnings (warning questionnaire)
# Questionnaire item Possible Answers
@@ 2 When you saw the warning dialogue, what was your first reaction? [free text]
3 I understood the warning dialogue. [5-point Likert scale, from “Strongly disagree” to “Strongly agree”]
4 I am familiar with this warning dialogue. [5-point Likert scale, from “Strongly disagree” to “Strongly agree”]
5 I am not interested in this warning dialogue. [5-point Likert scale, from “Strongly disagree” to “Strongly agree”]
@@ 6 Which word(s) did you find confusing or too technical? [free text]
7 Please rate the extent of risk you feel you were warned about. [very low risk; low risk; no risk; risky; very high risk]
8 What action, if any, did the warning dialogue want you to take? [to continue to the website; to be careful while continuing to the website; to not continue to the website; I did not feel anything]

### v2

@@ 3 Approach
Regarding the explanation style, the first approach we considered is based on feature-based explanations, a widely adopted approach in HCI systems for communicating the most influential features contributing to an AI decision (Lai2023SurveyHumanAIDecisionMaking; Fok24FeatureBased). In our context, these explanations highlight specific elements of an email, such as a masked link or a suspicious sender address, that led the AI to classify it as phishing. Complementing this standard, we also include counterfactual explanations, which offer an alternative, theoretically grounded approach to supporting user understanding. Rather than simply highlighting influential features, counterfactuals offer contrastive explanations answering implicit “why not?” questions by describing how small, meaningful changes would alter the AI’s prediction (Lee23Counterfactual; Koh24AlgorithmicRecourse; Gentile25Counterfactual). This explanation technique clarifies decision boundaries and supports actionable recourse, particularly relevant in high-stakes settings such as fraud detection or security. In our phishing warning dialogues, we embed these explanations to show how the email could have avoided being flagged, for example, stating that the email would be considered safe if the link label matched the actual destination URL (e.g., “protect your account” linking to facebook.com/protect-account).
The second factor concerns the LLM model used to generate the explanation message. While past studies often relied on manually written explanation messages to demonstrate their effectiveness (desolda2023explanations; greco2025enhancing; buono2023warnings), the growing availability of LLMs may enable the automatic generation of tailored content in both explanation styles. If explanations produced by LLMs prove as effective as manually crafted messages, which are both time-consuming to create and must be updated whenever new features are introduced, then LLMs could fully automate their generation. This study considered explanations generated by an open-source model (Llama 3.3) and a closed-source commercial model (Claude 3.5 Sonnet). These two LLMs were considered to evaluate whether a commercially available LLM (Claude) could provide better explanations than an open-weight solution such as LLama 3.3. This aspect is critical, as an organization may have strict privacy requirements that prohibit the sharing of sensitive data (e.g., emails) with third parties, such as Claude (Shashidhar2023LLMsOpenSource; Panwar2024ChatgptLlama; Zhang2024LLMsMedicalOpenCommercial; Adams2024LlamaProprietaryRadiology; Espejel2025LLMs).
@@
  • RQ4: How do user-related factors affect click-through rate (CTR) of phishing links?
@@ 3.1 Warning dialogues design
The phishing warning dialogues explored in this study have the same structure to isolate the effects of the explanation message, depicted in Figure 1. A dialogue opens with an attention-grabbing header flagging the link as potentially dangerous, followed by a summary that attributes this judgment to the AI system. The main part of the interface is the explanation message, presented as a short paragraph embedded within the dialogue box. The bottom part contains buttons to get back to the email visualization or, with a less prominent presentation, a link to continue browsing the destination. All messages adhere to the same typographic style and layout to minimize visual biases. This design ensured that any differences in user response could be attributed to the nature of the explanations rather than variations in formatting.
Figure 1: Anatomy of the warning dialogue used in the study. The top part (A) includes a brief warning that serves as the header of the box. The main part (B) presents the explanation message, while the footer (C) displays controls to return to the email visualization or proceed with browsing.
@@ 3.2 Explanation Styles
The explanation message included in the main part differs according to the considered styles. Feature-based explanations emphasize observable cues extracted from the email, such as sender identity mismatches or suspicious links. The left part of Figure 4 shows sample explanations for this style, realized through declarative statements that point to concrete phishing indicators. All messages share a common structure consisting of three sentences, inspired by prior work desolda2023explanations, which builds on the Communication-Human Information Processing (C-HIP) model (wogalter2018communication) and previous studies on effective warning communication (bauer2013warning). The same structure was explicitly adopted when prompting the LLMs to generate the warnings, ensuring consistency across human- and AI-authored messages.
In contrast, counterfactual explanations present a hypothetical scenario in which the email would not have been flagged. These explanations reframe the AI’s decision as contingent on a minor change, inviting users to consider alternative options. The examples adopt a conditional format, underscoring a key difference in framing: while feature-based explanations justify the current decision, counterfactuals highlight what could have made the outcome different. The right part of Figure 4 shows sample warnings using a counterfactual style.
@@ 3.3 Generating the explanations
The other difference in the experimental conditions is the LLM used to generate the warning text. We selected the two LLMs to create explanations for this study through a rigorous evaluation process. Notably, 13 different LLMs (six commercially available and seven open-source) were included in our evaluation process as potential candidates. As grounds for this evaluation, we considered the three emails (two true positives and one false positive) whose links, when clicked, generated a warning dialogue during the experiment. To evaluate each LLM, we followed this process:
  • Following the approach and prompts proposed in desolda2025apollo, for each of the three emails, the LLM was prompted to generate an explanation message to be included in the warning, given the email body and one of the features to be explained. We considered the following three email features, which were found to be helpful to users when making decisions regarding phishing content (greco2025enhancing; desolda2025apollo), i.e., (i) the Top-Level Domain of the URL in the email is misplaced (i.e., “www.amazon.com.cz”); (ii) the URL in the email is an IP address; (iii) the link shown in the email and its actual destination mismatch (e.g., ”click here” is shown instead of ”www.facebook.com/…”) — this feature was found in the false positive email. The full prompts, extracted from desolda2025apollo, are reported in Appendix 9.
  • For each of the three emails, we generated a counterfactual explanation message through a modified version of the prompt proposed in desolda2025apollo. This prompt is also reported in Appendix 9.
@@
  • The explanations underwent an automatic measurement of text readability with the FK (Flesch-Kincaid) Grade Level (kincaid1975derivation) and of text understandability with the SMOG (Simple Measure of Gobbledygook) Index (McLaughlin1969SMOG).
@@
To select the best commercial and open-weight LLMs, we initially selected a set of LLMs with the lowest, thus best, SMOG Index and FK Grade for both feature-based and counterfactual explanations. From this analysis, we selected Claude 3.5 Sonnet and Gemini 1.5 Flash as the best candidates for commercial LLMs, while all the Llama LLMs are candidates for the open-weight solutions. Then, two experts in HCI and cybersecurity performed a manual inspection of the generated explanations. This analysis removed Gemini 1.5 Flash because it generated explanations that were either too generic or contained hallucinations, especially for the counterfactual style (e.g., for the Amazon phishing email, it generated an explanation reporting “[..] A safe email address wouldn’t have strange parts in its name. [..]”). Moreover, Llama 3.2 90B (as well as Gemini 1.5 Pro) was excluded, as it could not be forced to produce an explanation for the false-positive email. Llama 3.1 8B was a good candidate for feature-based explanations, although readability and understandability metrics were mediocre in the context of counterfactual explanations. Moreover, the generated feature-based explanation for the Amazon email contained a hallucination (“[..] It looks like an Amazon link, but the ’.com’ name is typically found in other types of companies. [..]”). Therefore, considering the qualitative assessment and the average metrics for both explanation styles, we selected Llama 3.3 70B as it represents a good trade-off between the resources needed to deploy this LLM and the quality of the explanations. The explanations produced by Claude 3.5 Sonnet and Llama 3.3 70B were then featured in the user study as described in Section 4.1.
It is worth noting that while our companion work APOLLO (desolda2025apollo) validates the technical scalability of generating explanations for thousands of emails, this study adopts a Human-in-the-Loop validation approach for the experimental stimuli. We manually verified the LLM outputs, not because the models cannot generate text at scale, but to ensure that the experimental stimuli presented to the N=750 participants were free of hallucinations that could invalidate the internal validity of the user study. Thus, we validate the potential of high-quality automated explanations.
Figure 2: Average understandability and readability metrics for the feature-based explanations generated by the different LLMs — the lower, the better. The asterisks indicate the LLMs that were finally chosen.
Figure 3: Average understandability and readability metrics for the counterfactual explanations generated by the different LLMs — the lower, the better. The asterisks indicate the LLMs that were finally chosen.
@@ 3.4 Claude vs Llama
Instead, Llama delivers explanations that are somewhat more concise and technical in tone. Its feature-based outputs highlight specific anomalies (e.g., numeric IP addresses, suspicious domain extensions) with minimal elaboration. The counterfactuals follow a similar structure to Claude’s but are generally more direct and less descriptive, focusing on the change required to alter the classification without extending into consequences. Llama also tends to use simpler language and more neutral phrasing, which may be more mechanical but still effective. These stylistic differences are illustrated in Figure 4. The top row includes examples generated with Claude 3.5 Sonnet, while the bottom row includes examples generated with Llama 3.3 70B.
Figure 4: Four versions of the warning dialogue for the same deceptive link. Rows correspond to the two LLMs used in this study (Claude 3.5 Sonnet and Llama 3.3 70B), while columns correspond to explanation styles (feature-based and counterfactual).
@@ 4 Methods
In this section, we present a controlled experiment that compares the effectiveness of various explanations displayed in warning dialogues to protect users from phishing emails. The study specifically focused on evaluating explanations based on two factors we identified and discussed in the previous section: the explanation style (either feature-based or counterfactual) and the LLM model used to generate the explanation message (LLama 3.3 or Claude 3.5 Sonnet). We also considered a baseline warning that included manually generated messages, which was empirically demonstrated to outperform all existing warnings for phishing attacks (greco2025enhancing). The study procedure and materials (i.e., warning design and emails included) were identical to those in greco2025enhancing to ensure a fair and rigorous comparison with the baseline.
@@ 4.1 Study Design
This study employed a between-subjects design, with the explanation style and LLM used to generate the explanations as the independent variables. Five total between-subject levels were considered, four experimental warnings plus a baseline:
@@
    Baseline: warnings that included feature-based explanations, which were manually written by experts (existing dataset - from greco2025enhancing)
To improve the external validity of the study, for each experimental condition, we generated three explanations related to three different features. The complete list of explanation messages contained in each warning is reported in Appendix 10.
We excluded a ”no-explanation” as a baseline since previous research has explicitly established its inferiority. Notably, greco2025enhancing demonstrated that feature-based explanations significantly reduce phishing CTR (from 20.8% to 11.8%, $p=.032$) compared to generic warnings. Similarly, buono2023warnings validated the superiority of active, explanatory warnings over passive indicators. Consequently, our study focuses on benchmarking automated LLM generation against this expert-crafted ”gold standard”, rather than re-validating the known baseline efficacy.
@@ 4.2 Participants
A total of 750 participants were involved, with equal allocation (n = 150) to each condition. The manually generated feature-based explanation condition included existing data (collected separately, n = 150), while the remaining 600 participants were recruited specifically for this study through the Prolific online platform [^2] https://www.prolific.co. The inclusion criteria for participants were as follows: i) fluency in English, ii) an even split between men and women (50-50), and iii) 18 or more years of age. A total of 61 participants either failed attention checks or were identified as inattentive and were thus excluded. To balance participation across the experimental conditions, we recruited an additional 61 participants to reach a final sample of 750 participants. The final sample had an average age of 32.98 years (sd = 11.01) and reported spending 7.94 hours per day on the Internet (sd = 3.95). The participants consisted of 375 males and 375 females. The participation in the study lasted approximately 20 minutes, and participants were rewarded with £3.00, in line with Prolific’s recommended participation fee of £9.00/hour.
@@ 4.3 Materials
We purposely employed a limited set of phishing scenarios (2 phishing emails out of 14) to maintain a realistic prevalence ratio ( 14%) without extending the session duration beyond 20 minutes, which would induce fatigue effects. While this limits the generalizability across all possible phishing vectors, it ensures high ecological validity regarding user attention spans in a simulated work environment. Future work should adopt longitudinal designs to test a broader array of threats.
A warning dialogue, such as that shown in Figure 1, was shown to users who clicked on a phishing link. To simulate a case in which the system mistakenly classifies a genuine email as phishing, i.e., a false positive situation, we forced the platform to display a warning for one of the genuine emails as well. After the study, each user may have seen a maximum of three warning dialogues.
A full suite of user interactions was recorded in this study, including hovering over phishing links, clicking on them, visiting associated URLs, and responding to warnings, which enabled the calculation of the click-through rate (CTR) for each participant. The CTR was the proportion of accessed suspicious links to the number of displayed warnings, serving as an objective measure of the warning’s effectiveness.
The survey instrument employed by desolda2023explanations was administered to comprehend users’ perceptions of the warnings. All the questions and their formats of responses are given in Table 1.
To assess the effects of users’ knowledge of basic cybersecurity concepts, a 10-item test questionnaire was employed (olmstead2017securityquestionnaire).
Finally, a short, 6-item version of the Need for Cognition Scale (NCS-6) (coelho2020needforcognition6item) was used to measure users’ Need for Cognition, which is the tendency to engage in and enjoy effortful cognitive activities (cacioppo1982needforcognition). Previous work highlighted that individual differences like NFC might significantly impact the effectiveness of AI support and explanations in terms of performance, overreliance on AI, and learning (Cacioppo1996NFCCognitive; Bucinca2021NFC; Gajos2022NFC; Bucinca2024Optimizing)
Table 1: Questionnaire used to evaluate warning dialogues
# Question Possible answers
@@ 2 When you saw the warning dialogue, what was your first reaction? [free text]
3 I understood the warning dialogue. [5-point Likert: Strongly disagree – Strongly agree]
4 I am familiar with this warning dialogue. [5-point Likert: Strongly disagree – Strongly agree]
5 I am not interested in this warning dialogue. [5-point Likert: Strongly disagree–Strongly agree]
@@ 6 Which word(s) did you find confusing or too technical? [free text]
7 Please rate the extent of risk you feel you were warned about. [very low; low; no risk; risky; very high]
8 What action, if any, did the warning dialogue want you to take? [continue; be careful while continuing; do not continue; nothing]